Calculate the mass of Mg needed to produce 80 mL of hydrogen gas at STP.


1
Expert's answer
2021-04-06T06:07:43-0400

Mg + 2H2O "\to" Mg(OH)2 + H2

80 mL (0.08L) of Hydrigen at STP

moles of H2 produces = 0.08/22.4 = 0.00357 mol

mol of Mg required = 0.00357 mol

Mass of Mg = 0.00357×24 = 0.085 g